OpenStudy (anonymous):

1/A+1/B=5 Solve for A PLEASE HELP FINAL EXAM TOMORROW WILL MEDAL AND FAN

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

B+A=5AB

OpenStudy (skullpatrol):

why?

OpenStudy (anonymous):

Make the denominator the same and cancel

OpenStudy (skullpatrol):

Correct :-)

OpenStudy (anonymous):

but what do i do after that?

OpenStudy (skullpatrol):

Get A alone on one side of the equation, that's what "solve for A" means.

OpenStudy (anonymous):

but subtract B so cancel out B?

OpenStudy (skullpatrol):

B+A=5AB B = 5AB - A B = (5B - 1)A B/(5B - 1) = A <---solved for A

OpenStudy (anonymous):

where did the 1 come from?

OpenStudy (anonymous):

and I'm confused..why did you subtract A and didnt subtract B

OpenStudy (skullpatrol):

The idea is to get all the A's on one side and the B's on the other, then factor.

OpenStudy (anonymous):

but B is in there too? sorry im really confused

OpenStudy (skullpatrol):

B+A=5AB B = 5AB - A <---factor out A B = (5B - 1)A B/(5B - 1) = A <---solved for A

OpenStudy (anonymous):

ohhhh okay i see now. thank you!!

OpenStudy (skullpatrol):

thanks for asking :)